Show your solution.




1. Two charges of π‘ž1=+ 12.0 𝑛𝐢 and π‘ž2=βˆ’ 12.0 𝑛𝐢 are 0.100 m apart (such pairs of point charges with equal magnitude and opposite sign are called electric dipoles.). Calculate the electric field caused by π‘ž1.




2. Two charges of π‘ž1=+ 3.0 𝑛𝐢 and π‘ž2=βˆ’ 7.0 𝑛𝐢 are separated by a distance of 50 cm. What is the electric field strength at a point that is 20cm from π‘ž1 and 30cm from π‘ž2? The point lies between π‘ž1 and π‘ž2.





1
Expert's answer
2022-04-11T16:55:37-0400

1.

"E=\frac{kq_1}{r^2}=10.8~\frac{kN}C."

2.

"E=k(\frac{q_1}{r_1^2}-\frac{q_2}{r_2^2})=1375~\frac NC."